MazeRunner’s net income is reported as $1,300,000. The firm’s depreciation is $130,000. Assume the capital spending is $286,000 and the working capital of the firm increases by $91,000. What is the market value of equity if the FCFE is projected to grow at 5.7% indefinitely and the cost of equity is 17%? (Do not round intermediate calculations.)
Answer #1

FCFE = Net Income + Depreciation - Capex - Change in Net working Capital + New debt - Debt Repayment

= 1300000 + 130000 - 286000 - 91000 = $1,053,000

Market Value of equity = FCFE1/(k-g) = $1,053,000*1.057/(0.17-0.057) = $9.85 million